OpenStudy (anonymous):

HELP PLEASE Find the perimeter of the following shape, rounded to the nearest tenth: Shape ABCD is shown. Point A is at negative 3, 5. Point B is at 2, 6. Point C is at 0, 2. Point D is at negative 5, 1. 17.9 18 18.6 19.1

OpenStudy (anonymous):

@YinYangInsane @zepdrix

OpenStudy (ankitshaw):

do you know how to find the distance between two points?

OpenStudy (khalilforthewin):

wht grade

OpenStudy (anonymous):

no, and 10th.

OpenStudy (khalilforthewin):

fan me ill answer them all in a second

OpenStudy (anonymous):

19.1

OpenStudy (anonymous):

 distance AB^2 =(1^2+5^)=1+25=26 so AB =sq.rt.26 distance AD^2=(2^2+4^2)=4+16=20 so AD =sq.rt.20 perimeter =2(AB+AD) =19.14